# SaToR-G Collaboration

### v1

@@ Abstract
Figure 4: Behavior of $\alpha_{1}$ (black continuos line) as the phase of the demodulation sinusoid varies. The frequency of the sinusoid is fixed at the annual value. The dashed line represents the corresponding values for the standard deviation. The filter parameters are the same as those used in Figure 3.
The value of $\alpha_{1}$ at the phase $\phi=\phi_{0}\simeq 0.91$ rad (black circle) corresponds to the specific case of Figure 3. It is worth emphasizing that the maximum value of $\alpha_{1}$ is not obtained at the phase $\phi=\phi_{0}$, but at a higher value equal to approximately 2.18 rad. This reinforces our conclusion that what we obtain at the expected frequency and phase value for the violation, actually represents a null measurement for the PPN parameter.
We know of a number of perturbative effects with annual periodicity are present in the residuals of Figure 1. We have therefore estimated the phase of the solid and ocean tides and the phase related to the coefficient of the quadrupole moment of the Earth’s gravitational field [13]. The latter, like the tides, is characterized by a periodic oscillation (superimposed on a secular trend) with a primary component with annual periodicity, and is responsible for the largest systematic error in the measurement of the longitude of the LAGEOS II satellite. For the phase expected for the possible violation, the contribution of the main systematic errors of gravitational origin is about $1.5\times 10^{-5}$.
So even the analysis of the systematic errors, although it does not fully explain the residuals which have an annual periodicity, leads us overall to the conclusion of a measurement consistent with zero for $\alpha_{1}$.
Conclusions— LLI represents one of the cornerstones of both the standard model of field and particle physics and the standard model of gravitation, i.e. GR. In a sense, LLI represents our current deepest understanding of the nature of space and time. So, why test LLI? A strong motivation in our work is to search for the possible evidence of new physics beyond GR. We mentioned the possible existence of additional fields that come into play in mediating the gravitational interaction and that could couple to matter in such a way that, in some cases, they violate Lorentz Invariance. In this work we have presented and discussed a test of LLI in the gravitational sector by searching the possible existence of PFE.
We obtained a new constraint on the PPN parameter $\alpha_{1}$ from the analysis of the orbit of the geodetic satellite LAGEOS II. The result achieved is $\alpha_{1}=(3\pm 3)\times 10^{-5}$. The analysis has been performed on a time interval of approximately 28.3 years of the satellite orbit exploiting the precise measurements provided by the SLR technique and the main models of the GEODYN II orbit determination program.
This measurement is obtained exploiting a phase sensitive detection of the rate of the orbital residuals of LAGEOS II argument of pericenter and mean anomaly. The estimation of the main sources of systematic errors is dominated by the uncertainty on the Earth’s gravitational field quadrupole coefficient $\bar{C}_{2,0}$, about $2\times 10^{-5}$ and of the same order of magnitude of the standard deviation estimated from the PSD analysis. We highlighted how the measured value for $\alpha_{1}$ can correlate with the contribution of systematic errors of a gravitational nature when the contribution of the phase that we estimated for these effects at the annual frequency is introduced into the PSD analysis. The result is therefore compatible with a null value for the parameter $\alpha_{1}$, in agreement with the prediction of GR.
This is the first measurement that constraints the PPN parameter $\alpha_{1}$ from the satellite orbital data in the terrestrial field. The measurement improves the previous constraint for $\alpha_{1}$ obtained by means of the LLR technique [10] in the field of the Sun by a factor of 3. The result we have obtained further constrains the possible existence of a preferred frame for local gravitational physics and, consequently, that of theories of gravitation described, in addition to the metric tensor of GR, by the presence of additional fields of tensor and/or vector nature, such as for example the case of Einstein-aether theory [41], i.e. of vector-tensor theories of gravitation  [^10] Very interesting, the Einstein-aether theory can be seen as the low-energy limit of the Hořava-Lifshitz quantum theory [46]..
@@
Acknowledgments— This work has been performed by the SaToR-G (Satellite Test of Relativistic Gravity) collaboration funded by the Commissione Scientifica Nazionale (CSN2) of the Istituto Nazionale di Fisica Nucleare (INFN), to which we are very grateful. The authors acknowledge the ILRS for providing high-quality laser ranging data of the LAGEOS II satellite. This work is partially supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.

### v2

@@ Abstract
Figure 4: Behavior of $\alpha_{1}$ (solid line: blue-GEODYN II and red-SATAN) as the phase of the demodulation sinusoid varies. The frequency of the sinusoid is fixed at the annual value. The blue and red dashed lines represent the corresponding values for the standard deviation. The two circles, blue and red, correspond to the values of $\alpha_{1}$ obtained for GEODYN II and SATAN respectively, see Eq. (12). The green continuous line shows the overall estimated systematic error to $\alpha_{1}$ due to both gravitational and non-gravitational perturbations, about $8\times 10^{-6}$. The filter parameters are the same as those used in Figure 3.
The values of $\alpha_{1}$ at the phases $\phi=\phi_{0}\simeq 0.91$ rad (blue circle) and $\phi=\phi_{0}\simeq 2.56$ rad (red circle) correspond to the specific cases of Figure 3. Both in the case of the GEODYN II and SATAN analyses, it is worth emphasizing that the maximum values of $\alpha_{1}$ are not obtained at the phase $\phi=\phi_{0}$, but at a higher value equal to approximately 2.18 rad for GEODYN II, and at a lower value around 1.79 rad for SATAN. This reinforces our conclusion that what we obtain at the expected frequency and phase value for the violation, actually represents a null measurement for the PPN parameter. In the figure, the horizontal green line represents the overall systematic error $\lesssim 8\times 10^{-6}$ in the measurement of $\alpha_{1}$, which we consider to be independent of the phase of the possible violation signal. A detailed study of the main systematic errors is reported in the companion paper [15]. The budget is dominated by the uncertainty in the Earth’s hexadecapole coefficient $\bar{C}_{4,0}$, which contributes an error of $\sim 6.5\times 10^{-6}$. To assess the impact of non-gravitational perturbations (NGPs), we performed a rigorous end-to-end simulation, processing our high-fidelity NGP models through the entire measurement pipeline. This analysis confirms that the systematic bias from NGPs is negligible, with the largest contribution from SRP model errors being $(-0.4\pm 4.0)\times 10^{-6}$. The detailed error budget described in [15] leads to three significant conclusions: i) it underscores the thoroughness of our non-gravitational force modeling, ii) our analysis is demonstrably limited by statistics and not by systematics, and iii) the statistical error is due to broadband noise around the annual period. Hence, the analysis of the systematic errors reinforces the conclusion of a measurement consistent with zero for $\alpha_{1}$  [^9] The estimate of $\alpha_{1}$ is robust against systematic effects. Varying key parameters, such as those of the dynamical model and the Solar System’s velocity relative to the CMB, induces changes down to $\sim 10^{-7}$, well below the statistical uncertainty..
Conclusions— Local Lorentz Invariance (LLI) is a cornerstone of both the Standard Model of particle physics and General Relativity, and as such, it underpins our fundamental understanding of spacetime. Testing LLI is therefore a powerful probe for new physics. Many alternative theories of gravity introduce additional fields that can mediate the gravitational interaction, potentially violating LLI and giving rise to preferred-frame effects (PFE). In this Letter, we present a new test of LLI in the gravitational sector by searching for PFE in the orbits of the LAGEOS and LAGEOS II satellites. Our analysis spans nearly three decades of precise Satellite Laser Ranging data and employs two independent orbit determination software for cross-validation. The constraint is derived from a phase-sensitive detection of an annual periodic effect in the orbital residuals of the satellites’ mean argument of latitude. Using the POD software SATAN, we found $\alpha_{1}=(+1\pm 2)\times 10^{-5}$ for the Parameterized Post-Newtonian (PPN) parameter, while with GEODYN II we obtained a consistent result of $\alpha_{1}=(+2\pm 3)\times 10^{-5}$. The total systematic error budget is dominated by the uncertainty in the Earth’s hexadecapole harmonic $\bar{C}_{4,0}$ and is estimated to be $\sigma_{\text{sys}}\approx 8\times 10^{-6}$, which is about 2 times smaller than the statistical uncertainty obtained with SATAN. This result is consistent with a null value for $\alpha_{1}$, in full agreement with the predictions of General Relativity. The robustness of this null result was further confirmed through extensive testing, detailed in [15]. An analysis of the data using a shorter lock-in integration time (500 days) produced a time series of nearly uncorrelated $\alpha_{1}$ estimates that shows no significant temporal variation, confirming the stability of our result. Furthermore, an extreme sensitivity analysis demonstrated that our measurement is insensitive to large, pessimistic variations in key model parameters. This work provides the first constraint on the PPN parameter $\alpha_{1}$ using satellite orbital data in the Earth’s gravitational field. It improves upon the previous best constraint from Lunar Laser Ranging [10] by more than a factor of four, placing stringent new limits on the existence of a preferred frame for local gravitational physics and, consequently, on theories that feature additional vector or tensor fields, such as Einstein-aether theory [44] [^10] We note that Einstein-aether theory can be regarded as the low-energy limit of Hořava-Lifshitz gravity [49].. In [15] we give the bounds on the parameter space and coupling constants of Einstein’s-aether theory and a detailed comparison of our result for $\alpha_{1}$ with other tests of Lorentz invariance.
@@
Acknowledgments— This work has been performed by the SaToR-G (Satellite Test of Relativistic Gravity) collaboration funded by the Commissione Scientifica Nazionale (CSN2) of the Istituto Nazionale di Fisica Nucleare (INFN), to which we are very grateful. The authors acknowledge the International Laser Ranging Service (ILRS) for providing high-quality laser ranging data of the LAGEOS and LAGEOS II satellites and an anonymous Referee for helpful comments and remarks that have considerably improved the original Letter. This work is partially supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU.